40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 6 01:09:11 2017, johannes, Update, Computer Scripts / Programs, New set of loss measurements XARM_20170905.pdfYARM_20170905.pdf
    Reply  Mon Sep 11 12:56:40 2017, johannes, Update, Computer Scripts / Programs, lossmap attempts 
       Reply  Mon Sep 11 16:46:09 2017, Steve, Update, PEM, earthquakes M5.4eq.png
          Reply  Tue Sep 19 14:00:41 2017, Steve, Update, PEM, earthquakes 7.1MX5.3J.png
             Reply  Fri Oct 27 09:34:20 2017, Steve, Update, PEM, earthquakes recentEQs.png
                Reply  Fri Nov 10 13:46:26 2017, Steve, Update, PEM, M3.1 local earthquake 3.1M_local_EQ.png
                   Reply  Mon Apr 9 08:39:39 2018, Steve, Update, PEM, M5.3 eq Souther CA M5.3_Santa_Cruz_Is.CA.pngafter_M5.3.pngM5.3vac.png
                Reply  Thu Jan 25 08:27:25 2018, Steve, Update, PEM, earthquakes M4_local_eq.pngM4.pngM5.8Ferndale_CA.png
       Reply  Mon Sep 25 17:56:33 2017, johannes, Update, Computer Scripts / Programs, transmitted power during lossmap 
    Reply  Thu Oct 11 13:40:51 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
       Reply  Fri Oct 12 12:29:34 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
          Reply  Fri Oct 12 20:20:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
             Reply  Sat Oct 13 20:11:10 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                Reply  Mon Oct 15 10:32:13 2018, yuki, Update, Computer Scripts / Programs, loss measurements XARM_20181015_1500.pdfYARM_20181015_1500.pdf
                   Reply  Tue Oct 16 00:44:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                   Reply  Fri Nov 2 19:25:16 2018, gautam, Update, Computer Scripts / Programs, loss measurements 
    Reply  Mon Oct 15 12:52:54 2018, yuki, Update, Computer Scripts / Programs, additional comments 
       Reply  Fri Nov 2 16:42:31 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
          Reply  Mon Nov 5 13:52:18 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
             Reply  Tue Nov 6 10:19:26 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
                Reply  Tue Nov 6 15:23:48 2018, gautam, Update, IOO, IMC problematic MCwonky.png
                   Reply  Tue Nov 6 19:02:35 2018, aaron, Update, IOO, IMC problematic 181106_MCTRANS.jpg
                      Reply  Sat Nov 10 17:40:00 2018, aaron, Update, IOO, IMC problematic 
                         Reply  Mon Nov 12 13:53:20 2018, rana, Update, IOO, loss measurement: oscope vs CDS DAQ 
                         Reply  Thu Nov 15 10:21:07 2018, aaron, Update, IOO, IMC problematic 181115_chansDown.pngPD_noise.png
                            Reply  Fri Nov 16 10:53:07 2018, aaron, Update, IOO, IMC problematic Screenshot_from_2018-11-16_19-22-34.png
                               Reply  Sat Nov 17 18:59:01 2018, aaron, Update, IOO, IMC problematic 
                Reply  Wed Nov 7 05:16:16 2018, yuki, Update, Computer Scripts / Programs, arm loss measuremenents 
Message ID: 14270     Entry time: Mon Nov 5 13:52:18 2018     In reply to: 14268     Reply to this: 14274
Author: aaron 
Type: Update 
Category: Computer Scripts / Programs 
Subject: arm loss measuremenents 

After running this script Friday night, i noticed Saturday that the data hadn't saved. Scrolling up inthe terminal, I couldn't see where I'd run the script, so I thought I'd forgotten to run it as I was making last minute changes to the scope settings Friday before leaving.

Monday it turns out I hadn't forgotten to run the script, but the script itself was getting hung up as it waited for ASS to settle, due to the offset on the ETM PIT or YAW setpoints. The script was waiting until both pitch and yaw settled to below 0.7, but yaw was reading ~15; I think this is normal, and it looks like Yuki had solved this problem by waiting for the DEMOD-OFFSET to become small, rather than just the DEMOD signal to be small. Since this is a solved problem, I think I might be using an old script, but I'm pretty sure I'm running the one in Johannes' folder that Yuki is referencing for example here. The scripts in /yutaro_scripts/ have this DEMOD-OFFSET functionality commented out, and anyway those scripts seem to do the 2D loss maps rather than 1D loss measurements.

In the meantime I blocked the beams and ran the script in DARK mode. The script is saving data in /armloss/data/run_20181105/, and runs with no exceptions thrown.

However, when I try to dither align the YARM, I get an error that "this is not a degree of freedom that has an ASS". I'm alsogetting some exceptions from MEDM about unavailable channels. It must have been something about donatella not initializing, because it's working on pianosa. I turned on YARM ADS from pianosa. Monitoring from dataviewer, I see that LSC-TRY_OUT has some spikes to 0.5, but it's mostly staying near 0. I tried returning to the previous frozen outputs, and also stepping around ETMY-[PIT/YAW] from the IFO_ALIGN screen, but didn't see much change in the behavior of LSC-TRY. I missed the other controls Gautam was using to lock before, and I've also made myself unclear on whether ASS is acting only on angular dof, or also on length.

I unblocked the beams after the DARK run was done.

ELOG V3.1.3-